OpenStudy (anonymous):

Simplify 1 / (1+a^n) + 1/(1+a^-n)

OpenStudy (freckles):

hint: multiply that second fraction by a^n/a^n

OpenStudy (anonymous):

Thanks but does it cancel out th denominater for the second fraction

OpenStudy (freckles):

\[1 \cdot a_n=? \\ (1+a^{-n}) \cdot a_n=1 \cdot a_n +a^{-n} \cdot a^{n} =?\]

OpenStudy (anonymous):

Wouldn't it be 1+a as the denimontar

OpenStudy (freckles):

those one n's were suppose to be exponents (not subscipts)

OpenStudy (freckles):

\[1 \cdot a^n=a^n \\ (1+a^{-n}) \cdot a^n=1 \cdot a^n+a^{-n} a^{n} \text{ by distributive law } \\ \text{ now do you know law of exponents? }\]

OpenStudy (freckles):

if you have the same base and you are multiplying what do you do with the exponents ?

OpenStudy (anonymous):

Add

OpenStudy (freckles):

\[(1+a^{-n})a^n=a^n+a^{-n+n}=?\]

OpenStudy (freckles):

-n+n=?

OpenStudy (anonymous):

0

OpenStudy (freckles):

right and a^0=?

OpenStudy (anonymous):

One

OpenStudy (freckles):

\[\frac{1}{1+a^{n}}+\frac{a^n}{1+a^{n}}=?\]

OpenStudy (freckles):

you see you have the same denominator

OpenStudy (freckles):

now you can write as one fraction

OpenStudy (freckles):

\[\frac{1+a^{n}}{1+a^{n}}=?\]

OpenStudy (anonymous):

1

OpenStudy (freckles):

right and this is of coursing assuming a>0

OpenStudy (freckles):

a could be less than 0 depending on n we could say a lot about the domain restrictions lol

OpenStudy (freckles):

but I'm sure they are just looking for 1

OpenStudy (anonymous):

Ok tysm

OpenStudy (freckles):

np